# X-Portrait: Expressive Portrait Animation with Hierarchical Motion Attention

### v1

@@ Abstract.
stable diffusion, ControlNet, generative model, portrait animation, talking head
journal: TOGccs: Computing methodologies  Animations
1. Introduction
@@
Commencing with the pioneering works (Siarohin et al., 2019b, a), portrait animation has predominantly been approached as a two-step generative process (Zhao and Zhang, 2022; Siarohin et al., 2019b; Wang et al., 2021; Drobyshev et al., 2022), involving image warping and rendering. Specifically, the identity features, encoded from the given source image, undergo an initial warping step with the motion flow calculated based on the disparities in expressions and head poses between the source and driving frame. Subsequently, these warped source features are fed into a generative decoder to synthesize the final animated frames with inpainted facial details and background. Despite substantial advancement in synthesizing high-fidelity talking heads, such methods exhibit limitations in both warping and rendering processes. The warping field, generated from landmarks (Geng et al., 2018), neural keypoints (Wang et al., 2021) or latent code (Wang et al., 2022), often fails to capture subtle or extreme out-of-domain facial expressions. Additionally, due to the restricted capability of the decoder, the synthesized images are often limited in resolution and perceptual quality. Unnatural artifacts, such as blurriness and rigid shading variation, and facial distortions become apparent when animating out-of-domain portraits or driving with long-range head motions.
@@
With the recent advent of text-to-image diffusion models (Ho et al., 2020; Song et al., 2020a, b; Rombach et al., 2022), we have witnessed unprecedented diversity and stability in image synthesis demonstrated by large diffusion models pre-trained on billions of images, such as Imagen (Saharia et al., 2022) and Stable Diffusion (SD) (AI, 2022). Consequently we aim to harness the generative capability of production-ready diffusion models, specifically SD1.5 in our work, for the task of portrait animation. Recently a line of work (Hu et al., 2023; Chang et al., 2024; Xu et al., 2023b; Zhang et al., 2023b) has approached this task as a problem of controlled image-to-video (I2V) diffusion. In particular, the appearance context from the source image is cross-queried by the self-attention layers of SD UNets, whereas the motion is controlled using explicitly derived semantic motion signals such as landmarks (Chang et al., 2024), skeleton (Hu et al., 2023; Chang et al., 2024) and dense pose (Xu et al., 2023b), within the framework of ControlNet (Zhang et al., 2023a). However, such explicit and coarse pose control signals do not fully convey the original expressions, and heavily rely on the robustness and accuracy of third-party pose detectors, hindering animation expressiveness and stability. Moreover, the controlling pose maps, derived from the driving frames, are not fully agnostic to the driving facial semantics (e.g., shape and ratio) and noticeably impact the source characteristics during cross-identity animation, referred to as appearance leakage here.
@@
In this work, we propose X-Portrait, a novel zero-shot framework that leverages image diffusion priors for expressive portrait animation, showcasing superior perceptual quality, motion richness, identity preservation, and domain generalization (Fig. X-Portrait: Expressive Portrait Animation with Hierarchical Motion Attention). Building upon prior controlled image-to-video works (Hu et al., 2023; Chang et al., 2024; Xu et al., 2023b), we innovate the pose control scheme to mitigate the aforementioned issues, namely expressiveness loss and appearance leakage. To fully retain the driving head poses and facial expressions, we argue that the network should interpret the motion directly from the original driving images, without resorting to any intermediate motion representation. Subsequently, we design our pose ControlNet (Zhang et al., 2023a) using the RGB images as conditional inputs. However, standard motion transfer is typically trained on videos in a self-driven scheme where the source and driving share the same identity. To avoid merely copying the appearance and structure directly from the conditional pose image by the model, we employ a refined state-of-the-art motion transfer network (Wang et al., 2021) to generate cross-identity training image pairs. Our cross driven training scheme simultaneously mitigates appearance leakage, enabling direct portrait animation during inference without any pre-processing. To further enhance the derivation of subtle facial expressions at nuanced scales, we deploy an auxiliary ControlNet (Zhang et al., 2023a) to guide the conditional motion attention to local facial movements especially around the eyes and mouth.
@@
  • A novel implicit motion control scheme trained with cross-identity scaling-augmented images, effectively retaining driving motion context to the fullest extent and simultaneously mitigating appearance leakage during inference.
@@
Figure 2. Overview of X-Portrait . For the task of portrait animation, X-Portrait  leverages a frozen pre-trained LDM as a rendering backbone, and incorporates three auxiliary trainable modules for disentangled control of appearance $\mathcal{R}$, motion $\mathcal{C}$ and temporal smoothness $\mathcal{M}$. Specifically, $\mathcal{R}$ extracts the source appearance and background context from a reference image $I_{S},$ and $\mathcal{C}$ derives the motion of head pose and facial expression from a driving frame $I_{D}.$ During training, we leverage a pre-trained network $\mathcal{F}$ to generate cross-identity control images $I_{C}$ as conditional input to our control modules $\mathcal{C}.$ To better capture subtle expressions, we enhance the attention to the local detailed facial movements with an additional masked control image $I_{C}^{l}.$ Both $I_{C}$ and $I_{C}^{l}$ are subject to random heterogeneous scaling for mitigation of appearance leakage from the drivings. For inference, we animate a source portrait directly with the video frames without any pre-processing, enabling expressive and robust animation with strictly maintained identity resemblance.
2. Related Work
2.1. GAN-based Portrait Animation
@@
A majority of methods for portrait animation leverage Generative Adversarial Networks (GANs) to learn motion dynamics in self-supervised ways. The pioneering works(Siarohin et al., 2019a, b) consists of two stages, i.e., warping and rendering. They utilize sparse neural keypoints to estimate head motion, with which the encoded source identity features are warped and fed into a generative decoder to synthesize animated frames with inpainted facial details and background. Following this framework, a line of approaches have focused on enhancing the representations for warping estimation, such as 3D neural keypoints (Face Vid2Vid (Wang et al., 2021)), depth(DaGAN (Hong et al., 2022)) and thin-plate splines (TPS  (Siarohin et al., 2021)). Additionally, ReenactArtFace (Qu et al., 2023) employed a 3D morphable model to rig the expressions and poses, whereas ToonTalker (Gong et al., 2023) leveraged a transformer architecture to facilitate the warping process when dealing with cross-domain datasets. To enhance rendered image quality, MegaPortraits (Drobyshev et al., 2022) advances the generator to megapixel resolution by leveraging high-resolution image data. FADM (Zeng et al., 2023) proposes a coarse-to-fine animation framework, employing a diffusion process to enrich facial details over the initial results generated by FOMM (Siarohin et al., 2019b) and Face Vid2Vid (Wang et al., 2021). In addition to video reenactment, there have also been advancements embracing diverse driving signals, such as 3D facial prior (Deng et al., 2020; Khakhulin et al., 2022; Sun et al., 2023; Feng et al., 2021; Xu et al., 2023a) and audios (Guo et al., 2021; Sun et al., 2022). Nevertheless, these methods are reliant on explicit source features warping, with a primary focus on talking scenarios. This paper aims to generate high-fidelity expressive animations over portraits with diverse domain styles.
2.2. Diffusion-based Portrait Animation
@@
Recently Diffusion Probabilistic Models (DMs) (Ho et al., 2020; Song et al., 2020a, b) have achieved superior performance in various generative tasks, such as image (Saharia et al., 2022), video (Blattmann et al., 2023), and multi-view renderings (Liu et al., 2023a, b; Gu et al., 2023). Latent diffusion models (Rombach et al., 2021) have further advanced this domain with reduced computational costs by operating the diffusion step in a lower-dimensional latent space. Relevant to our task of portrait animation, pre-trained diffusion models (Saharia et al., 2022; Rombach et al., 2021) have been leveraged in many image to video (I2V) works. Notably, several studies (Cao et al., 2023; Lin et al., 2023; Zhang, 2023) highlight the effectiveness of fusing features of a reference image into the self-attention blocks in the LDM UNets, facilitating image editing and video generation with preserved appearance context. Additionally, the nominal work ControlNet (Zhang et al., 2023a) extends the LDM framework to controllable image generation with additive structural conditions from signals such as landmarks, segmentations and dense poses. With those, several concurrent works (Hu et al., 2023; Xu et al., 2023b; Chang et al., 2024) achieve state-of-the-art full-body portrait animations by seamlessly integrating the appearance and motion control as well as temporal attentions (Guo et al., 2023b, a) with pre-trained UNets. However, all of their motion control signals, i,e, skeleton with/without facial landmarks  (Hu et al., 2023; Chang et al., 2024) and dense pose (Xu et al., 2023b), may not fully capture the original motion and depend heavily on the accuracy of third-party detectors, potentially hindering the animation expressiveness and stability. In contrast, we innovate the control signals to enable driving the portraits directly with a video, without resorting to any intermediate explicit motion representations.
3. Method
@@
Given as few as a single static portrait $I_{S}$, the objective of our approach is to generate a head animation sequence $\{I_{S\rightarrow D_{i}}\}$ with a length of $q$, conditioned on a driving video ${I_{D_{i}}}$, where $i=0,\ldots,q$ denotes the frame index. The generated image $I_{S\rightarrow D_{i}}$ is intended to preserve the identity characteristics and background content depicted in $I_{S}$ while accurately following the head pose and facial expressions featured in the driving frame $I_{D_{i}}$. Conventional portrait animation algorithms, such as Face Vid2Vid (Wang et al., 2021) (denoted as $\mathcal{F}$), are typically trained as a frame reconstruction task over a large corpus of videos, where $I_{S}$ and $I_{D}$ are both sourced from the same subject. During inference, the reference and driving images could feature distinct identities to achieve motion transfer.
@@
In this work, X-Portrait  leverages the powerful generative prior of production-ready Latent Diffusion Models, with disentangled control of appearance and temporal motion (Zhang et al., 2023a) (Section 3.1). On top of it, we propose a novel cross-identity training scheme with the aid of $\mathcal{F}$ that enables direct and accurate driving motion derivation from $I_{D}$ (Section 3.2). We further refine the control granularity by guiding the motion attention to local driving image patches with an auxiliary ControlNet (Zhang et al., 2023a) as elaborated in Section 3.3. Lastly, we provide details on several training and inference techniques for enhancing the retention of identity features from $I_{S}$ (Section 3.4). Figure. 2 provides an overview for the model architecture with the training and inference pipeline.
3.1. Preliminaries
@@
Latent Diffusion Model. The Diffusion Models (DM) (Ho et al., 2020; Song et al., 2020a, b) are generative models designed to synthesize desired data samples from Gaussian noise $z_{T}\sim N(0,1)$ through $T$ denoising steps. Latent diffusion models (Rombach et al., 2022) are a class of diffusion models that operates in the latent space facilitated by a pretrained auto-encoder. We note that, unlike standard Text-to-Image (T2I) models conditioned on textual inputs, our application derives scene context and motion constraints from $I_{S}$ and $I_{D}$ respectively without text description. Consequently, text control is intentionally excluded from our formulations. The model is then trained to learn the reverse denoising process with the objective,
(1) $L_{ldm}=\mathbb{E}_{z_{0},t,\epsilon\sim\mathcal{N}(0,1)}\bigg{[}\Big{\lVert}\epsilon-\epsilon_{\theta}\big{(}z_{t},t\big{)}\Big{\lVert}_{2}^{2}\bigg{]},$
@@
Portrait Reenactment with Diffusion. A line of works (Hu et al., 2023; Chang et al., 2024; Xu et al., 2023b) have recently delved into the utilization of pre-trained LDMs for full-body human video reenactment. While exhibiting slight algorithmic variations, these frameworks generally share a similar structure, consisting of plug-and-play modules in conjunction with a frozen Stable Diffusion (SD) UNet. Notably, these modules encompass: 1) an Appearance Reference Module $\mathcal{R}$ that extracts the identity attributes and background context from $I_{S}$ to be cross-queried by the self-attention blocks of the backbone UNets; 2) a Control Module $\mathcal{C}$, typically in a form of ControlNet (Zhang et al., 2023a), that establishes structural mappings between control signals $I_{C}$ and the output with additive attentions; 3) a Temporal Module $\mathcal{M}$ (Guo et al., 2023b),intervened within UNet blocks with temporal transformers, that ensures cross-frame correspondence for maintaining temporal coherence. We follow the same framework structure the but innovate on the controls, focusing on meticulous transfer of the source identity attributes as well as driving facial expressions and head poses.
@@
Figure 3. Ablation. (a) Training with scaling-augmented ground-truth driving image (with the same identity as reference) as motion condition results in severe identity appearance leakage from the driving. (b) Our local motion control module effectively enhances the capture of subtle and detailed local facial movements. (c) Random scaling with the conditional control images improves the identity preservation (note the head shape and eye sizes).
3.2. Implicit Cross-Identity Motion Control
@@
To achieve control of facial expressions and head poses with image diffusion models, prior works typically employ a ControlNet (Zhang et al., 2023a) trained to condition image generation on facial landmarks, as showcased in e.g., (Chang et al., 2024). The control module $\mathcal{C}$ is trained to reconstruct $I_{D}$ conditioned on the landmarks input extracted from the target $I_{D},$ with $I_{S}$ as the input to the appearance reference module $\mathcal{R}.$ $I_{S}$ and $I_{D}$ are two random video frames during training, featuring the same subject. While effective at a coarse scale, such a control scheme induces several problems (Figure 5), particularly when zoomed in on faces, as required in our task. Firstly, the accuracy of the driving signals is heavily dependent on the precision of third-party detectors, such as OpenPose (Cao et al., 2019). This dependence introduces jittered controls, motion ambiguity, and could result in corrupted animation when the detection fails, for example, due to face occlusion. Secondly, the conveyance of strong emotions or subtle expressions often involves detailed facial movements, such as those in the teeth, eyeballs, eyebrows, and ajna. The animation expressiveness is significantly hindered by the coarse landmark representation, which may not capture the nuances demanded for accurate facial animation. Lastly, the driving landmarks are aligned with the face structure of targeted image $I_{D},$ featuring the same subject as in $I_{S}$. Therefore under the self-driven training scheme, the network, as a short-cut, tends to copy the driving structure entangled with identity features such as facial shapes and ratios. As a result, undesirable identity drift to the driving subject occurs during cross-identity animation in inference.
@@
To address the aforementioned issues, we demand a novel conditional motion control that is entirely disentangled from the source identity features, while minimizing the loss of motion information at all scales, such as facial expressions and head poses. This motivates us to utilize the original driving RGB image $I_{D}$, featuring a different subject than $I_{S}$, as our conditional input. This aligns with our real-world applications as well, enabling the direct reenactment of the source image onto the driving video of a different identity. However, such image pairs with distinct identities but with aligned motions are not readily accessible for training.
@@
Facilitated by a pre-trained portrait reenactment network $\mathcal{F}$ such as Face Vid2Vid (Wang et al., 2021) in our work, we propose to synthesize cross-identity image pairs for the training of our control module $\mathcal{C}$. Specifically, two randomly selected video frames featuring the same subject are used, serving as the appearance reference $I_{S}$ and the reconstruction target $I_{D}$ respectively. However, instead of relying on facial landmarks from $I_{D}$, we employ $\mathcal{F}$ to generate an RGB control image $I_{C}=\mathcal{F}(I_{S^{\prime}}\rightarrow I_{D})$ as the conditional input to $\mathcal{C},$ where $I_{S^{\prime}}$ is a frame randomly selected from a video with a distinct identity. In contrast to explicit extracted motion control signals, our cross-identity training scheme effectively instructs the control module $\mathcal{C}$ to implicitly derive the identity-disentangled motion from $I_{C}$. That being said, this mitigates appearance leakage from the driving signal, allowing direct application of the driving video for inference without third-party dependency. Notably, while degenerates with extreme expressions and long-range head motion, $\mathcal{F}$ offers reenacted control image $I_{C}$ of reasonable quality and motion accuracy for widely-distributed conversational scenarios. Even with limited perceptual quality, our control signal $I_{C}$ contains richer motion information than landmarks, which is sufficient for $\mathcal{C}$ to decipher the embedded motion structure effectively, enabling it to adapt and correlate to finer expressions and poses when provided with ground-truth motions for supervision. Therefore, as illustrated in our experiments, our trained control module is able to establish implicit structural mapping between $I_{C}$ and $I_{D},$ generalizing well to unseen expressions and head motions.
3.3. Enhanced Attention to Local Facial Movements

### v3

@@ Abstract
{strip}[Uncaptioned image]
Figure 1: Given a single reference portrait (left column), X-Portrait  is capable of synthesizing compelling and expressive animations (right columns), covering large head pose changes and highly dynamic and detailed facial expression features from the input driving videos. Notably, X-Portrait  can faithfully preserve the identity information from the reference portrait while transferring the expression subtleties across a wide range of portrait styles. Please see the supplementary video for more dynamic results.
1 Introduction
@@
Commencing with the pioneering works [35, 34], portrait animation has predominantly been approached as a two-step generative process [51, 35, 43, 10], involving image warping and rendering. Specifically, the identity features, encoded from the given source image, undergo an initial warping step with the motion flow calculated based on the disparities in expressions and head poses between the source and driving frame. Subsequently, these warped source features are fed into a generative decoder to synthesize the final animated frames with inpainted facial details and background. Despite substantial advancement in synthesizing high-fidelity talking heads, such methods exhibit limitations in both warping and rendering processes. The warping field, generated from landmarks [12], neural keypoints [43] or latent code [44], often fails to capture subtle or extreme out-of-domain facial expressions. Additionally, due to the restricted capability of the decoder, the synthesized images are often limited in resolution and perceptual quality. Unnatural artifacts, such as blurriness and rigid shading variation, and facial distortions become apparent when animating out-of-domain portraits or driving with long-range head motions.
@@
With the recent advent of text-to-image diffusion models [18, 37, 38, 32], we have witnessed unprecedented diversity and stability in image synthesis demonstrated by large diffusion models pre-trained on billions of images, such as Imagen [33] and Stable Diffusion (SD) [1]. Consequently we aim to harness the generative capability of production-ready diffusion models, specifically SD1.5 in our work, for the task of portrait animation. Recently a line of work [21, 6, 46, 48] has approached this task as a problem of controlled image-to-video (I2V) diffusion. In particular, the appearance context from the source image is cross-queried by the self-attention layers of SD UNets, whereas the motion is controlled using explicitly derived semantic motion signals such as landmarks [6], skeleton [21, 6] and dense pose [46], within the framework of ControlNet [50]. However, such explicit and coarse pose control signals do not fully convey the original expressions, and heavily rely on the robustness and accuracy of third-party pose detectors, hindering animation expressiveness and stability. Moreover, the controlling pose maps, derived from the driving frames, are not fully agnostic to the driving facial semantics (e.g., shape and ratio) and noticeably impact the source characteristics during cross-identity animation, referred to as appearance leakage here.
@@
In this work, we propose X-Portrait, a novel zero-shot framework that leverages image diffusion priors for expressive portrait animation, showcasing superior perceptual quality, motion richness, identity preservation, and domain generalization (Fig. 1). Building upon prior controlled image-to-video works [21, 6, 46], we innovate the pose control scheme to mitigate the aforementioned issues, namely expressiveness loss and appearance leakage. To fully retain the driving head poses and facial expressions, we argue that the network should interpret the motion directly from the original driving images, without resorting to any intermediate motion representation. Subsequently, we design our pose ControlNet [50] using the RGB images as conditional inputs. However, standard motion transfer is typically trained on videos in a self-driven scheme where the source and driving share the same identity. To avoid merely copying the appearance and structure directly from the conditional pose image by the model, we employ a refined state-of-the-art motion transfer network [43] to generate cross-identity training image pairs. Our cross driven training scheme simultaneously mitigates appearance leakage, enabling direct portrait animation during inference without any pre-processing. To further enhance the derivation of subtle facial expressions at nuanced scales, we deploy an auxiliary ControlNet [50] to guide the conditional motion attention to local facial movements especially around the eyes and mouth.
@@
  • A novel implicit motion control scheme trained with cross-identity scaling-augmented images, effectively retaining driving motion context to the fullest extent and simultaneously mitigating appearance leakage.
@@
Figure 2: Overview of X-Portrait . For the task of portrait animation, X-Portrait  leverages a frozen pre-trained LDM as a rendering backbone, and incorporates three auxiliary trainable modules for disentangled control of appearance $\mathcal{R}$, motion $\mathcal{C}$ and temporal smoothness $\mathcal{M}$. Specifically, $\mathcal{R}$ extracts the source appearance and background context from a reference image $I_{S},$ and $\mathcal{C}$ derives the motion of head pose and facial expression from a driving frame $I_{D}.$ During training, we leverage a pre-trained network $\mathcal{F}$ to generate cross-identity control images $I_{C}$ as conditional input to our control modules $\mathcal{C}.$ To better capture subtle expressions, we enhance the attention to the local detailed facial movements with an additional masked control image $I_{C}^{l}.$ Both $I_{C}$ and $I_{C}^{l}$ are subject to random heterogeneous scaling for mitigation of appearance leakage from the drivings. For inference, we animate a source portrait directly with the video frames without any pre-processing, enabling expressive and robust animation with strictly maintained identity resemblance.
2 Related Work
2.1 GAN-based Portrait Animation
@@
A majority of methods for portrait animation leverage Generative Adversarial Networks (GANs) to learn motion dynamics in self-supervised ways. The pioneering works[34, 35] consists of two stages, i.e., warping and rendering. They utilize sparse neural keypoints to estimate head motion, with which the encoded source identity features are warped and fed into a generative decoder to synthesize animated frames with inpainted facial details and background. Following this framework, a line of approaches have focused on enhancing the representations for warping estimation, such as 3D neural keypoints (Face Vid2Vid [43]), depth(DaGAN [20]) and thin-plate splines (TPS  [36]). Additionally, ReenactArtFace [30] employed a 3D morphable model to rig the expressions and poses, whereas ToonTalker [13] leveraged a transformer architecture to facilitate the warping process when dealing with cross-domain datasets. To enhance rendered image quality, MegaPortraits [10] advances the generator to megapixel resolution by leveraging high-resolution image data. FADM [47] proposes a coarse-to-fine animation framework, employing a diffusion process to enrich facial details over the initial results generated by FOMM [35] and Face Vid2Vid [43]. In addition to video reenactment, there have also been advancements embracing diverse driving signals, such as 3D facial prior [8, 23, 40, 11, 45] and audios [15, 41]. Nevertheless, these methods are reliant on explicit source features warping, with a primary focus on talking scenarios. This paper aims to generate high-fidelity expressive animations over portraits with diverse domain styles.
2.2 Diffusion-based Portrait Animation
@@
Recently Diffusion Probabilistic Models (DMs) [18, 37, 38] have achieved superior performance in various generative tasks, such as image [33], video [3], and multi-view renderings [26, 25, 14]. Latent diffusion models [31] have further advanced this domain with reduced computational costs by operating the diffusion step in a lower-dimensional latent space. Relevant to our task of portrait animation, pre-trained diffusion models [33, 31] have been leveraged in many image to video (I2V) works. Notably, several studies [4, 24, 49] highlight the effectiveness of fusing features of a reference image into the self-attention blocks in the LDM UNets, facilitating image editing and video generation with preserved appearance context. Additionally, the nominal work ControlNet [50] extends the LDM framework to controllable image generation with additive structural conditions from signals such as landmarks, segmentations and dense poses. With those, several concurrent works [21, 46, 6] achieve state-of-the-art full-body portrait animations by seamlessly integrating the appearance and motion control as well as temporal attentions [17, 16] with pre-trained UNets. However, all of their motion control signals, i,e, skeleton with/without facial landmarks  [21, 6] and dense pose [46], may not fully capture the original motion and depend heavily on the accuracy of third-party detectors, potentially hindering the animation expressiveness and stability. In contrast, we innovate the control signals to enable driving the portraits directly with a video, without resorting to any intermediate explicit motion representations.
3 Method
@@
Given as few as a single static portrait $I_{S}$, the objective of our approach is to generate a head animation sequence $\{I_{S\rightarrow D_{i}}\}$ with a length of $q$, conditioned on a driving video ${I_{D_{i}}}$, where $i=0,\ldots,q$ denotes the frame index. The generated image $I_{S\rightarrow D_{i}}$ is intended to preserve the identity characteristics and background content depicted in $I_{S}$ while accurately following the head pose and facial expressions featured in the driving frame $I_{D_{i}}$. Conventional portrait animation algorithms, such as Face Vid2Vid [43] (denoted as $\mathcal{F}$), are typically trained as a frame reconstruction task over a large corpus of videos, where $I_{S}$ and $I_{D}$ are both sourced from the same subject. During inference, the reference and driving images could feature distinct identities to achieve motion transfer.
@@
In this work, X-Portrait  leverages the powerful generative prior of production-ready Latent Diffusion Models, with disentangled control of appearance and temporal motion [50] (Section 3.1). On top of it, we propose a novel cross-identity training scheme with the aid of $\mathcal{F}$ that enables direct and accurate driving motion derivation from $I_{D}$ (Section 3.2). We further refine the control granularity by guiding the motion attention to local driving image patches with an auxiliary ControlNet [50] as elaborated in Section 3.3. Lastly, we provide details on several training and inference techniques for enhancing the retention of identity features from $I_{S}$ (Section 3.4). Figure. 2 provides an overview for the model architecture with the training and inference pipeline.
3.1 Preliminaries
@@
Latent Diffusion Model. The Diffusion Models (DM) [18, 37, 38] are generative models designed to synthesize desired data samples from Gaussian noise $z_{T}\sim N(0,1)$ through $T$ denoising steps. Latent diffusion models [32] are a class of diffusion models that operates in the latent space facilitated by a pretrained auto-encoder. We note that, unlike standard Text-to-Image (T2I) models conditioned on textual inputs, our application derives scene context and motion constraints from $I_{S}$ and $I_{D}$ respectively without text description. Consequently, text control is intentionally excluded from our formulations. The model is then trained to learn the reverse denoising process with the objective,
$L_{ldm}=\mathbb{E}_{z_{0},t,\epsilon\sim\mathcal{N}(0,1)}\bigg{[}\Big{\lVert}\epsilon-\epsilon_{\theta}\big{(}z_{t},t\big{)}\Big{\lVert}_{2}^{2}\bigg{]},$ (1)
@@
Portrait Reenactment with Diffusion. A line of works [21, 6, 46] have recently delved into the utilization of pre-trained LDMs for full-body human video reenactment. While exhibiting slight algorithmic variations, these frameworks generally share a similar structure, consisting of plug-and-play modules in conjunction with a frozen Stable Diffusion (SD) UNet. Notably, these modules encompass: 1) an Appearance Reference Module $\mathcal{R}$ that extracts the identity attributes and background context from $I_{S}$ to be cross-queried by the self-attention blocks of the backbone UNets; 2) a Control Module $\mathcal{C}$, typically in a form of ControlNet [50], that establishes structural mappings between control signals $I_{C}$ and the output with additive attentions; 3) a Temporal Module $\mathcal{M}$ [17],intervened within UNet blocks with temporal transformers, that ensures cross-frame correspondence for maintaining temporal coherence. We follow the same framework structure the but innovate on the controls, focusing on meticulous transfer of the source identity attributes as well as driving facial expressions and head poses.
@@
Figure 3: Ablation. (a) Training with scaling-augmented ground-truth driving image (with the same identity as reference) as motion condition results in severe identity appearance leakage from the driving. (b) Our local motion control module effectively enhances the capture of subtle and detailed local facial movements. (c) Random scaling with the conditional control images improves the identity preservation (note the head shape and eye sizes).
3.2 Implicit Cross-Identity Motion Control
@@
To achieve control of facial expressions and head poses with image diffusion models, prior works typically employ a ControlNet [50] trained to condition image generation on facial landmarks, as showcased in e.g., [6]. The control module $\mathcal{C}$ is trained to reconstruct $I_{D}$ conditioned on the landmarks input extracted from the target $I_{D},$ with $I_{S}$ as the input to the appearance reference module $\mathcal{R}.$ $I_{S}$ and $I_{D}$ are two random video frames during training, featuring the same subject. While effective at a coarse scale, such a control scheme induces several problems (Figure 5), particularly when zoomed in on faces, as required in our task. Firstly, the accuracy of the driving signals is heavily dependent on the precision of third-party detectors, such as OpenPose [5]. This dependence introduces jittered controls, motion ambiguity, and could result in corrupted animation when the detection fails, for example, due to face occlusion. Secondly, the conveyance of strong emotions or subtle expressions often involves detailed facial movements, such as those in the teeth, eyeballs, eyebrows, and ajna. The animation expressiveness is significantly hindered by the coarse landmark representation, which may not capture the nuances demanded for accurate facial animation. Lastly, the driving landmarks are aligned with the face structure of targeted image $I_{D},$ featuring the same subject as in $I_{S}$. Therefore under the self-driven training scheme, the network, as a short-cut, tends to copy the driving structure entangled with identity features such as facial shapes and ratios. As a result, undesirable identity drift to the driving subject occurs during cross-identity animation in inference.
@@
To address the aforementioned issues, we demand a novel conditional motion control that is entirely disentangled from the source identity features, while minimizing the loss of motion information at all scales, such as expressions and head poses. This motivates us to utilize the original driving RGB image $I_{D}$, featuring a different subject than $I_{S}$, as our conditional input. This aligns with our real-world applications as well, enabling the direct reenactment of the source image onto the driving video of a different identity. However, such image pairs with distinct identities but with aligned motions are not readily accessible for training.
@@
Facilitated by a pre-trained portrait reenactment network $\mathcal{F}$ such as Face Vid2Vid [43] in our work, we propose to synthesize cross-identity image pairs for the training of our control module $\mathcal{C}$. Specifically, two randomly selected video frames featuring the same subject are used, serving as the appearance reference $I_{S}$ and the reconstruction target $I_{D}$ respectively. However, instead of relying on facial landmarks from $I_{D}$, we employ $\mathcal{F}$ to generate an RGB control image $I_{C}=\mathcal{F}(I_{S^{\prime}}\rightarrow I_{D})$ as the conditional input to $\mathcal{C},$ where $I_{S^{\prime}}$ is a frame randomly selected from a video with a distinct identity. In contrast to explicit extracted motion control signals, our cross-identity training scheme effectively instructs the control module $\mathcal{C}$ to implicitly derive the identity-disentangled motion from $I_{C}$. That being said, this mitigates appearance leakage from the driving signal, allowing direct application of the driving video for inference without third-party dependency. Notably, while degenerates with extreme expressions and long-range head motion, $\mathcal{F}$ offers reenacted control image $I_{C}$ of reasonable quality and motion accuracy for widely-distributed conversational scenarios. Even with limited perceptual quality, our control signal $I_{C}$ contains richer motion information than landmarks, which is sufficient for $\mathcal{C}$ to decipher the embedded motion structure effectively, enabling it to adapt and correlate to finer expressions and poses when provided with ground-truth motions for supervision. Therefore, as illustrated in our experiments, our trained control module is able to establish implicit structural mapping between $I_{C}$ and $I_{D},$ generalizing well to unseen expressions and head motions.
3.3 Enhanced Attention to Local Facial Movements